OpenStudy (anonymous):

If r(t)= 3t^2 i + 1t j + (t^2-10t) k gives the position of a particle at time , find the time at which the speed of the particle is minimized.

OpenStudy (anonymous):

looking for t=?

OpenStudy (zarkon):

I get t=1/2

OpenStudy (anonymous):

i need answer for the same question ..... looking for t= ?

OpenStudy (anonymous):

t =1/2 is not correct :(

OpenStudy (zarkon):

\[v(t)=[6t,1,2t-10]\]

OpenStudy (anonymous):

i need t=? not v(t)

OpenStudy (zarkon):

\[s(t)=\|v(t)\|=\sqrt{(6t)^2+1^2+(2t-10^2}\] \[=\sqrt{40t^2-40t+101}\]

OpenStudy (zarkon):

now you want to find the time for which the speed is a minimum.

OpenStudy (anonymous):

ok

OpenStudy (zarkon):

take \[\frac{d}{dt}s(t)\] set equal to zero and solve for t

OpenStudy (anonymous):

still im not getting the right answer what u get for t ?

OpenStudy (zarkon):

I get 1/2

OpenStudy (anonymous):

this question is from web work so it telling me is not corect

OpenStudy (zarkon):

is the problem typed correctly...is r(t) correct above?

OpenStudy (anonymous):

If r(t)= -(t^2)i+(8t)j+((t^2)-3t)k gives the position of a particle at time , find the time at which the speed

OpenStudy (anonymous):

thats my questions

OpenStudy (zarkon):

well then you should get a different answer

OpenStudy (anonymous):

ok can u solve for this one

OpenStudy (anonymous):

coz i did same way but still i m getting wrong

OpenStudy (zarkon):

3/4

OpenStudy (anonymous):

can u help me plz

OpenStudy (anonymous):

ya now is corect

OpenStudy (anonymous):

can u help me with diffrent question tooo

OpenStudy (zarkon):

if you do the procedure I did above with your problem you will get the 3/4

OpenStudy (anonymous):

ya now i get it i just miss _2

OpenStudy (anonymous):

i need help with this question

OpenStudy (anonymous):

Find the length of the curve given by r(t)=((2)^(1/2))*ti/2+e^(t/2))j+(e^(-t/2))k where -2<=t<=1

OpenStudy (zarkon):

just compute \[\int\limits_{-2}^{1}\sqrt{\left[\frac{dx(t)}{dt}\right]^2+\left[\frac{dy(t)}{dt}\right]^2+\left[\frac{dz(t)}{dt}\right]^2}dt\]

OpenStudy (anonymous):

ok then ?

OpenStudy (anonymous):

first part is 1/2

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

plz r u helping me coz i dont no how \[\int\limits_{?}^{?}\] this eq